CaCO3=CaO+CO2
suppose 83.0 L of csrbon dioxide gas are produced by this reaction at a temperature of 370.0 degrees celsius and pressure of exactly 1 atm. calculate the mads of calcium carbonate that must have reacted. Be sure your answer has the correct number of significant digits.
Answer #1

Here is the solution of your question. If you have any doubt or need any clarification please comment in comment box and will definitely resolve your query. If you find useful please upvote it. Thanks in advance.

a from PV = nRT Padatm Tz 370°C, = (273 +370) = 643 k. V=83.02 R=0.0821 L-am kimel By putting values n = Pu - latm x 83L. RT